# WavefrontDiffusion: Dynamic Decoding Schedule or Improved Reasoning

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have achieved remarkable progress in complex reasoning and structured generation tasks such as mathematical problem solving and code synthesis (openai2025gptoss120bgptoss20bmodel; deepseekai2025deepseekr1incentivizingreasoningcapability). Autoregressive (AR) models remain the dominant paradigm for these tasks due to their stepwise logical consistency (deletang2024language). However, their strictly sequential nature introduces latency and limits flexibility, which can be problematic in settings that demand both accuracy and responsiveness, such as interactive assistants or real-time code generation. These limitations have motivated the exploration of alternative decoding paradigms that can balance quality, efficiency, and adaptability (leviathan2023fastinferencetransformersspeculative).
Diffusion Language Models (DLMs) have recently emerged as a promising alternative by framing text generation as an iterative denoising process (gong2025diffucoderunderstandingimprovingmasked; song2025seeddiffusionlargescalediffusion). Unlike AR models, which finalize one token per step, DLMs update multiple tokens in parallel over a series of iterations (schiff2025simpleguidancemechanismsdiscrete; lou2024discrete). This design enables more flexible generation while maintaining global coherence, offering potential gains in both efficiency and representational capacity(sahoo2024simple).
The capabilities of DLMs partly rely on the design of their denoising schedule, yet this aspect has received limited exploration. Current mainstream strategies include Standard Diffusion(chang2022maskgit) and BlockDiffusion(arriola2025block). Standard Diffusion applies a global and uniform denoising policy without restricting the update range, which often causes overconfidence in end-of-sequence tokens(nie2025largelanguagediffusionmodels) and insufficient context for denoised tokens, leading to early errors that cannot be corrected. Block Diffusion improves stability by processing fixed-size blocks in a preset order, but it does not account for the natural semantic boundaries of the output, causing the denoising of coherent sequences to be artificially fragmented.
@@ 2.1 Preliminaries: Diffusion Language Models
Diffusion Language Models (DLMs) treat text generation as an iterative denoising process over a sequence of discrete tokens (austin2021structured). Let $\mathbf{x}_{0}=(x_{0,1},\dots,x_{0,N})$ denote a clean token sequence of length $N$. The forward corruption process gradually replaces tokens with a special mask token [MASK], following a masking probability $t\in[0,1]$:
@@ 4.1 Experimental Setup
Tasks and Datasets. We evaluate WavefrontDiffusion on four benchmarks that test both mathematical reasoning and program synthesis: GSM8K(cobbe2021trainingverifierssolvemath), MATH(saxton2019analysingmathematicalreasoningabilities), HumanEval(chen2021evaluating), and Big Bench Hard (BBH)(suzgun2022challenging). These datasets cover diverse domains and represent challenging evaluation scenarios for diffusion-based decoding.
All experiments are performed in a strict zero-shot setting without chain-of-thought prompting or few-shot examples. Any intermediate reasoning traces are generated by the model itself rather than injected hints. For reasoning benchmarks (GSM8K, MATH, BBH) we report accuracy using exact match (EM), and for HumanEval we report pass@1 accuracy. For semantic quality evaluation (Section 4.3), we use the WikiText dataset. We randomly sample 1,000 sentences and repeat evaluations several times, then report the averaged scores.
Model and Baselines. We use two discrete diffusion language models, LLaDA-8B-Instruct and LLaDA-1.5. Keeping the same backbone across all methods ensures that performance differences come only from the decoding schedule. We compare WavefrontDiffusion with two baselines: 1) Standard Diffusion, which updates all masked tokens in parallel at every step. This highlights the effect of structured scheduling. 2) BlockDiffusion, which denoises fixed, non-overlapping blocks in a rigid order and is the current state-of-the-art among block-based schedules.
@@ 4.2 Main Results
Table 1: Main Performance Comparison. Each dataset reports accuracy (Acc) and computation cost time (Cost). Highlighted rows indicate our proposed WavefrontDiffusion, which consistently outperforms BlockDiffusion under identical compute budgets.
Strategy GSM8K MATH HumanEval BBH
Acc Cost Acc Cost Acc Cost Acc Cost
@@
Standard Diffusion 23.15 110.1 26.60 108.8 17.68 115.8 11.30 122.7
BlockDiffusion 80.74 113.6 40.62 109.4 45.73 107.2 43.23 116.6
@@
Standard Diffusion 30.17 111.9 27.04 111.4 34.76 118.0 17.12 127.2
BlockDiffusion 82.33 111.6 41.64 107.4 46.34 110.1 44.56 118.6
@@
We first evaluate whether WavefrontDiffusion improves reasoning and code generation accuracy when the compute budget is fixed. Table 1 shows results across four benchmarks. WavefrontDiffusion achieves the best performance on all tasks and for both model scales. It outperforms Standard Diffusion by a large margin and also provides consistent gains over BlockDiffusion.
Analysis. For the larger LLaDA-8B-Instruct, WavefrontDiffusion improves accuracy by +1.27 on GSM8K, +0.42 on MATH, +1.83 on HumanEval, and +1.07 on BBH compared to BlockDiffusion. For the smaller LLaDA-1.5, the improvements are +0.61 on GSM8K, +0.32 on MATH, +1.83 on HumanEval, and +0.70 on BBH. These gains are obtained with the same number of forward steps (1024) and under the same runtime budget. Although the reported wall-clock cost may differ slightly due to measurement noise, the computational load is identical across methods.
The results highlight two key points. First, adaptive scheduling helps the model finalize tokens only when sufficient context is available. This reduces premature commitments that often occur in BlockDiffusion and improves answer correctness. Second, the improvements appear consistently across both mathematical reasoning and program synthesis, and across different model sizes, showing that the benefits of dynamic scheduling are general and not tied to scale. Together, these findings confirm that the performance gains come from better scheduling rather than additional compute.
@@ 4.3 Semantic Quality Evaluation
To address RQ2, we test whether WavefrontDiffusion produces outputs that are more faithful and coherent than BlockDiffusion. We use BERTScore (zhang2020bertscoreevaluatingtextgeneration), which compares system outputs with human references through contextual embeddings. Recall (R) reflects completeness of the output, Precision (P) reflects avoidance of irrelevant tokens, and F1 balances both.
@@ 5 Related Work
Diffusion Language Models Diffusion models originated in image generation (ho2020denoisingdiffusionprobabilisticmodels; song2020generativemodelingestimatinggradients). Recent research extends these principles to text, leading to Diffusion Language Models (DLMs) (li2022diffusionlmimprovescontrollabletext). DLMs generate text through an iterative denoising process, showing promise in various text generation tasks (ye2025dream7bdiffusionlarge; xie2025dreamcoder7bopendiffusion). Early DLMs employed global masking strategies, predicting all masked tokens in parallel at each denoising step(li2023diffusionmodelsnonautoregressivetext). DLMs offer advantages in long text generation and mitigating exposure bias (zeng2025treediffastguidedcodegeneration).
Decoding Strategies for DLM Decoding strategies in text generation optimize output quality. Autoregressive models commonly use greedy decoding, beam search, or sampling methods (meister2022bestfirstbeamsearch). These methods typically generate tokens sequentially. For non-autoregressive models, particularly DLMs, decoding strategies are crucial. Existing DLM decoding strategies include: Standard Diffusion: This method repredicts all masked tokens at each timestep, lacking exploitation of already finalized tokens. Block Diffusion: This strategy partitions the sequence into fixed blocks for parallel generation. However, it can disrupt semantic coherence due to rigid boundaries. These strategies often involve trade-offs between generation quality and computational efficiency.
Confidence-Based Generation Confidence-based generation methods leverage a model’s certainty in its predictions to guide the generation process. Such approaches are common in non-autoregressive models, enhancing both generation efficiency and quality (gulrajani2023likelihoodbaseddiffusionlanguagemodels; li2025surveydiffusionlanguagemodels). For instance, some methods select high-confidence tokens for finalization based on prediction certainty, focusing subsequent steps on lower-confidence regions. This strategy prioritizes stabilizing accurate predictions, thereby reducing error accumulation.

### v3

# arXiv:2511.19473v3[cs.LG] 02 Mar 2026

@@ 1 Introduction
Recent advances in large language models (LLMs) have achieved remarkable progress in complex reasoning and structured generation tasks such as mathematical problem solving and code synthesis (OpenAI et al., 2025; DeepSeek-AI et al., 2025). Autoregressive (AR) models remain the dominant paradigm for these tasks due to their stepwise logical consistency (Deletang et al., 2024). However, their strictly sequential nature introduces latency and limits flexibility, which can be problematic in settings that demand both accuracy and responsiveness, such as interactive assistants or real-time code generation. These limitations have motivated the exploration of alternative decoding paradigms that can balance quality, efficiency, and adaptability (Leviathan et al., 2023).
Diffusion Language Models (DLMs) have recently emerged as a promising alternative by framing text generation as an iterative denoising process (Gong et al., 2025; Song et al., 2025). Unlike AR models, which finalize one token per step, DLMs update multiple tokens in parallel over a series of iterations (Schiff et al., 2025; Lou et al., 2024). This design enables more flexible generation while maintaining global coherence, offering potential gains in both efficiency and representational capacity(Sahoo et al., 2024).
The capabilities of DLMs partly rely on the design of their denoising schedule, yet this aspect has received limited exploration. Current mainstream strategies include Standard Diffusion(Chang et al., 2022) and BlockDiffusion(Arriola et al., 2025). Standard Diffusion applies a global and uniform denoising policy without restricting the update range, which often causes overconfidence in end-of-sequence tokens(Nie et al., 2025) and insufficient context for denoised tokens, leading to early errors that cannot be corrected. Block Diffusion improves stability by processing fixed-size blocks in a preset order, but it does not account for the natural semantic boundaries of the output, causing the denoising of coherent sequences to be artificially fragmented.
@@ 2.1 Preliminaries: Diffusion Language Models
Diffusion Language Models (DLMs) treat text generation as an iterative denoising process over a sequence of discrete tokens (Austin et al., 2021a). Let $\mathbf{x}_{0}=(x_{0,1},\dots,x_{0,N})$ denote a clean token sequence of length $N$. The forward corruption process gradually replaces tokens with a special mask token [MASK], following a masking probability $t\in[0,1]$:
@@ 3.1 Method Theory
Theoretical Grounding. We formalize this intuition through the Information Gradient Hypothesis, which posits that the conditional entropy of a token increases monotonically with its distance from the finalized context. While BlockDiffusion restricts the search space to rigid pre-defined blocks, WavefrontDiffusion performs an optimal restricted search within ”entropy isosurfaces” defined by this distance. We prove in Appendix D that our dynamic boundary is mathematically guaranteed to contain a higher density of low-entropy candidates compared to static blocks, thereby minimizing the probability of semantic mismatch.
@@ 4.1 Experimental Setup
Tasks and Datasets. We evaluate WavefrontDiffusion on five benchmarks that test both mathematical reasoning and program synthesis: GSM8K(Cobbe et al., 2021), MATH(Saxton et al., 2019), HumanEval(Chen et al., 2021), MBPP(Austin et al., 2021b), and Big Bench Hard (BBH)(Suzgun et al., 2022). These datasets cover diverse domains and represent challenging evaluation scenarios for diffusion-based decoding.
All experiments are performed in a strict zero-shot setting without chain-of-thought prompting or few-shot examples. Any intermediate reasoning traces are generated by the model itself rather than injected hints. For reasoning benchmarks (GSM8K, MATH, BBH) we report accuracy using exact match (EM), and for HumanEval and MBPP we report pass@1 accuracy. For semantic quality evaluation (Section 4.3), we use the WikiText dataset. We randomly sample 1,000 sentences and repeat evaluations several times, then report the averaged scores.
Model and Baselines. We use three discrete diffusion language models: LLaDA-8B-Instruct, LLaDA-1.5, and Dream-7B (Ye et al., 2025). Keeping the same backbone across all methods ensures that performance differences come only from the decoding schedule. We compare WavefrontDiffusion with two baselines: 1) Standard Diffusion, which updates all masked tokens in parallel at every step. This highlights the effect of structured scheduling. 2) BlockDiffusion, which denoises fixed, non-overlapping blocks in a rigid order and is the current state-of-the-art among block-based schedules.
@@ 4.2 Main Results
Table 1: Main Performance Comparison. Each dataset reports accuracy (Acc) and computation cost time (Cost). Highlighted rows indicate our proposed WavefrontDiffusion, which consistently outperforms BlockDiffusion under identical compute budgets. MBPP and Dream-7B results are added to demonstrate generalization.
Strategy GSM8K MATH HumanEval MBPP BBH
Acc Cost Acc Cost Acc Cost Acc Cost Acc Cost
@@
Standard Diffusion 23.15 110.1 26.60 108.8 17.68 115.8 13.50 109.1 11.30 122.7
BlockDiffusion 80.74 113.6 40.62 109.4 45.73 107.2 41.17 109.2 43.23 116.6
@@
Standard Diffusion 30.17 111.9 27.04 111.4 34.76 118.0 17.04 114.4 17.12 127.2
BlockDiffusion 82.33 111.6 41.64 107.4 46.34 110.1 44.04 118.1 44.56 118.6
@@
Dream-7B
Standard Diffusion 35.03 99.8 29.98 100.8 20.12 99.5 25.05 99.1 16.27 103.2
BlockDiffusion 78.92 99.2 43.60 103.5 53.05 100.6 58.52 95.9 45.13 100.5
[…]
We first evaluate whether WavefrontDiffusion improves reasoning and code generation accuracy when the compute budget is fixed. Table 1 shows results across five benchmarks. WavefrontDiffusion achieves the best performance on all tasks and for all three model families. It outperforms Standard Diffusion by a large margin and also provides consistent gains over BlockDiffusion.
Analysis. For the larger LLaDA-8B-Instruct, WavefrontDiffusion improves accuracy by +1.27 on GSM8K, +0.42 on MATH, +1.83 on HumanEval, +1.23 on MBPP, and +1.07 on BBH compared to BlockDiffusion. For the smaller LLaDA-1.5, the improvements are +0.61 on GSM8K, +0.32 on MATH, +1.83 on HumanEval, +2.16 on MBPP, and +0.70 on BBH. We observe similar consistent gains on the Dream-7B model (e.g., +1.74 on GSM8K, +1.22 on HumanEval, and +0.51 on MBPP), confirming that the benefits of our method generalize across different architectures.
These gains are obtained with the same number of forward steps (1024) and under the same runtime budget. Although the reported wall-clock cost may differ slightly due to measurement noise, the computational load is identical across methods.
The results highlight two key points. First, adaptive scheduling helps the model finalize tokens only when sufficient context is available. This reduces premature commitments that often occur in BlockDiffusion and improves answer correctness. Second, the improvements appear consistently across both mathematical reasoning and program synthesis, and across different model sizes and families, showing that the benefits of dynamic scheduling are general and not tied to scale. Together, these findings confirm that the performance gains come from better scheduling rather than additional compute.
@@ 4.3 Semantic Quality Evaluation
To address RQ2, we test whether WavefrontDiffusion produces outputs that are more faithful and coherent than BlockDiffusion. We use BERTScore (Zhang et al., 2020), which compares system outputs with human references through contextual embeddings. Recall (R) reflects completeness of the output, Precision (P) reflects avoidance of irrelevant tokens, and F1 balances both.
@@ 5 Related Work
Diffusion Language Models Diffusion models originated in image generation (Ho et al., 2020; Song and Ermon, 2020). Recent research extends these principles to text, leading to Diffusion Language Models (DLMs) (Li et al., 2022). DLMs generate text through an iterative denoising process, showing promise in various text generation tasks (Ye et al., 2025; Xie et al., 2025). Early DLMs employed global masking strategies, predicting all masked tokens in parallel at each denoising step(Li et al., 2023). DLMs offer advantages in long text generation and mitigating exposure bias (Zeng et al., 2025).
Decoding Strategies for DLM Decoding strategies in text generation optimize output quality. Autoregressive models commonly use greedy decoding, beam search, or sampling methods (Meister et al., 2022). These methods typically generate tokens sequentially. For non-autoregressive models, particularly DLMs, decoding strategies are crucial. Existing DLM decoding strategies include: Standard Diffusion: This method repredicts all masked tokens at each timestep, lacking exploitation of already finalized tokens. Block Diffusion: This strategy partitions the sequence into fixed blocks for parallel generation. However, it can disrupt semantic coherence due to rigid boundaries. These strategies often involve trade-offs between generation quality and computational efficiency.
Confidence-Based Generation Confidence-based generation methods leverage a model’s certainty in its predictions to guide the generation process. Such approaches are common in non-autoregressive models, enhancing both generation efficiency and quality (Gulrajani and Hashimoto, 2023; Li et al., 2025). For instance, some methods select high-confidence tokens for finalization based on prediction certainty, focusing subsequent steps on lower-confidence regions. This strategy prioritizes stabilizing accurate predictions, thereby reducing error accumulation.
@@ 6 Conclusion
Acknowledgements
We thank the anonymous reviewers for their constructive comments and suggestions, which helped improve the clarity and quality of this work. The work is partially supported by the U.S. National Science Foundation (NSF) Grant CRII 2451683, an NVIDIA Academic Grants Program, a U.S. Bank Academic Research Award, the University of California, Merced, and a UC Merced Faculty Research Award. The views and conclusions are those of the authors and do not necessarily reflect the official policy or position of the U.S. Government.
@@ B.6 Detailed Accuracy of BBH
Appendix C Additional Experimental Results
C.1 Comparison with Advanced Dynamic Baselines
To further validate the effectiveness of WavefrontDiffusion, we extend our comparison to include three recent advanced decoding strategies that also incorporate dynamic or adaptive mechanisms. These methods represent the state-of-the-art in optimizing diffusion decoding beyond fixed schedules:
  • Truncated-BlockDiffusion (TBD) (Anonymous, 2025): A concurrent approach that adaptively adjusts the block size during generation based on local confidence, aiming to solve the rigidity of fixed blocks.
  • Running Confidence Remasking (RCR) (He et al., 2025): A training-free strategy proposed in MDPO that identifies and re-masks low-confidence tokens during the decoding process, allowing the model to revise potential errors.
  • Self-Speculative Decoding (SSD) (Gao et al., 2025): A method designed primarily for acceleration via hierarchical verification, which effectively acts as a dynamic acceptance policy for generated tokens.
Results. We conduct the comparison using LLaDA-8B-Instruct with the total step budget fixed at $T=1024$. As shown in Table 9, WavefrontDiffusion consistently outperforms these advanced baselines across all four benchmarks. While TBD and RCR improve upon Standard Diffusion, they still lag behind WavefrontDiffusion in complex reasoning tasks (e.g., MATH and GSM8K). This suggests that while adaptive block sizing (TBD) and remasking (RCR) are beneficial, the wavefront expansion strategy—which aligns update order with semantic dependencies—provides a more optimal decoding path for reasoning-heavy generation.
Table 9: Comparison with Advanced Dynamic Baselines. Accuracy on LLaDA-8B-Instruct ($T=1024$). WavefrontDiffusion achieves state-of-the-art performance among training-free dynamic strategies.
Method GSM8K MATH HumanEval BBH
Standard Diffusion 23.15 26.60 17.68 11.30
BlockDiffusion 80.74 40.62 45.73 43.23
Truncated-BlockDiffusion (TBD) 80.90 40.84 45.73 39.41
Running Confidence Remasking (RCR) 80.60 40.02 43.90 43.10
Self-Speculative Decoding (SSD) 79.15 38.88 44.51 41.25
[…]
C.2 Efficiency and Computational Parity Analysis
A key concern in iterative decoding is whether performance gains come at the cost of increased latency. We clarify that WavefrontDiffusion is a scheduling strategy, not an acceleration method, designed to align strictly with the computational budget of standard block-based decoding.
Computational Parity. As detailed in Table 10, our method shares the identical number of denoising steps ($T=1024$) and model configuration as BlockDiffusion. WavefrontDiffusion alters the order of token updates but not the quantity of forward passes. Consequently, the total FLOPs per sample remain equivalent ($2.517\times 10^{15}$ for LLaDA-8B). Observed wall-clock time differences are within $\pm 2\%$, attributed solely to system variance rather than algorithmic overhead.
Table 10: Compute-Parity Comparison. We report FLOPs, total inference time (s), and throughput (Tokens Per Second, TPS) on LLaDA-8B and LLaDA-1.5 ($T=1024$). WavefrontDiffusion maintains identical computational cost to BlockDiffusion while achieving higher accuracy.
Model Strategy FLOPs GSM8K MATH HumanEval BBH
($\times 10^{15}$) Time TPS Time TPS Time TPS Time TPS
LLaDA-8B Standard Diffusion 2.517 110.1 9.30 108.8 9.41 115.8 8.84 122.7 8.35
BlockDiffusion 2.517 113.6 9.02 109.4 9.36 107.2 9.55 116.6 8.78
WavefrontDiffusion 2.517 112.2 9.13 109.9 9.32 109.3 9.37 115.1 8.89
LLaDA-1.5 Standard Diffusion 2.517 111.9 9.15 111.4 9.19 118.0 8.68 127.2 8.05
BlockDiffusion 2.517 111.6 9.17 107.4 9.53 110.1 9.30 118.6 8.63
WavefrontDiffusion 2.517 109.3 9.37 110.9 9.23 112.9 9.07 114.1 8.97
Compatibility with Acceleration. Although WavefrontDiffusion itself does not aim to accelerate inference, it is fully compatible with existing acceleration frameworks. Specifically:
  • KV-Cache Optimization: Since our method only changes the masking schedule, it integrates seamlessly with memory-reuse techniques like Fast-dLLM (Wu et al., 2025).
  • Speculative Decoding: The adaptive wavefront expansion is orthogonal to sampling compression and hierarchical verification methods (Gao et al., 2025), allowing for combined use to achieve both quality improvements and latency reduction.
C.3 Ablation Study on Denoising Steps
To investigate how the number of denoising steps affects generation quality, we performed a controlled ablation on LLaDA-8B-Instruct using the GSM8K benchmark. We varied the total diffusion steps $T\in\{256,512,1024\}$, which correspond to finalize quotas $k_{t}\in\{4,2,1\}$ respectively.
As shown in Table 11, WavefrontDiffusion remains competitive even at lower denoising steps and consistently surpasses BlockDiffusion as $T$ increases. The relative improvement becomes more pronounced at higher $T$, indicating that the adaptive wavefront scheduling effectively leverages extended denoising horizons to propagate semantic dependencies.
Table 11: Ablation on Denoising Steps ($T$). Accuracy on GSM8K using LLaDA-8B-Instruct. WavefrontDiffusion scales effectively with denoising depth.
Strategy Steps = 256 Steps = 512 Steps = 1024
Standard Diffusion 12.96 17.05 23.15
BlockDiffusion 43.14 67.02 80.74
[…]
C.4 Robustness Analysis: Temperature and Random Seeds
We further evaluated the robustness of WavefrontDiffusion under stochastic sampling conditions. We conducted experiments on GSM8K (LLaDA-8B), varying the sampling temperature from 0.0 to 0.8. Each value represents the mean accuracy $\pm$ standard deviation over three random seeds.
Table 12 demonstrates that WavefrontDiffusion maintains superior performance and low variance across all temperature regimes. The consistent advantage over BlockDiffusion (approx. +1.6%–2.0%) indicates that our dynamic frontier scheduling is robust to sampling randomness.
Table 12: Temperature Robustness Analysis. Mean accuracy $\pm$ standard deviation on GSM8K (LLaDA-8B).
Strategy Temp = 0.0 Temp = 0.3 Temp = 0.6 Temp = 0.8
Standard Diffusion 23.15 $\pm$ 0.00 23.42 $\pm$ 0.15 22.97 $\pm$ 0.21 20.35 $\pm$ 0.22
BlockDiffusion 80.74 $\pm$ 0.00 80.83 $\pm$ 0.14 81.09 $\pm$ 0.20 78.42 $\pm$ 0.37
[…]
C.5 Effectiveness with Alternative Token-Priority Metrics
To verify that our method’s success is not dependent solely on confidence scores, we evaluated WavefrontDiffusion using two alternative frontier-priority metrics on the Dream-7B model:
  1. 1.
    Entropy: Lower entropy indicates higher priority.
  2. 2.
    Margin: Larger gap between top-1 and top-2 probabilities indicates higher priority.
As shown in Table 13, WavefrontDiffusion consistently outperforms baselines regardless of the priority metric used, confirming the universality of the wavefront scheduling mechanism.
Table 13: Performance under Alternative Token-Priority Metrics. Experiments on Dream-7B with equal FLOPs ($T=1024$).
Strategy Metric GSM8K MATH HumanEval BBH
Standard Diffusion Entropy 34.80 29.56 20.13 15.90
BlockDiffusion Entropy 81.96 42.26 50.61 44.63
[…]
Standard Diffusion Margin 34.90 29.28 20.73 16.10
BlockDiffusion Margin 78.85 41.16 50.00 44.90
[…]
C.6 Matched-Latency Comparison with Autoregressive Models
Finally, we provide a latency-matched comparison against Autoregressive (AR) greedy decoding and Self-Speculative Decoding (SSD). Table 14 shows that WavefrontDiffusion achieves comparable latency to AR decoding while delivering superior accuracy, demonstrating its practical viability.
Table 14: Matched-Latency Comparison. LLaDA-8B-Instruct on GSM8K.
Method Latency (ms/token) $\downarrow$ Accuracy $\uparrow$
Autoregressive Greedy 9.2 80.67
Self-Speculative Decoding 5.8 79.30
[…]
C.7 Calibration Analysis
Since WavefrontDiffusion relies on confidence scores to determine the decoding frontier, validating that these scores are well-calibrated is critical. We conducted a rigorous calibration analysis to ensure that the model’s internal confidence serves as a reliable signal for decision-making.
Experimental Setup. To measure intrinsic calibration without the confounding factor of divergent reasoning paths, we employed a standard Teacher Forcing protocol. We fed the model the ground truth prefix at each step and recorded the confidence of tokens selected by WavefrontDiffusion versus BlockDiffusion. We report the Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) on GSM8K, MATH, and HumanEval.
Quantitative Results. Table 15 summarizes the calibration metrics. WavefrontDiffusion achieves consistently lower ECE scores across all domains, with a reduction of approximately 40% on GSM8K ($0.1093\to 0.0659$). Furthermore, it significantly reduces the Maximum Calibration Error (MCE) on complex tasks like MATH and HumanEval ($0.7051\to 0.4502$). This indicates that our adaptive wavefront effectively filters out poorly calibrated tokens that static block schedules are forced to commit to.
Table 15: Calibration Metrics. Comparison of Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) under Teacher Forcing. Lower is better.
Dataset Method ECE ($\downarrow$) MCE ($\downarrow$)
GSM8K BlockDiffusion 0.1093 0.4727
WavefrontDiffusion (Ours) 0.0659 0.4824
MATH BlockDiffusion 0.1025 0.3457
WavefrontDiffusion (Ours) 0.0901 0.2393
HumanEval BlockDiffusion 0.0920 0.7051
WavefrontDiffusion (Ours) 0.0853 0.4502
Reliability Diagrams. Figure 3 visualizes the reliability diagrams for both methods across the three datasets. Ideally, the calibration curve should align with the diagonal (dashed line). Comparing the left column (BlockDiffusion) and the right column (WavefrontDiffusion), we observe that WavefrontDiffusion’s curves are consistently closer to the diagonal, particularly in high-confidence regions. This confirms that the tokens selected by the wavefront are fundamentally more ”trustworthy,” validating the reliability of our adaptive scheduling mechanism.
(a) GSM8K - BlockDiffusion
(b) GSM8K - WavefrontDiffusion
(c) MATH - BlockDiffusion
(d) MATH - WavefrontDiffusion
(e) HumanEval - BlockDiffusion
(f) HumanEval - WavefrontDiffusion
Figure 3: Reliability Diagrams Comparison. Calibration curves for BlockDiffusion (Left) and WavefrontDiffusion (Right) across GSM8K, MATH, and HumanEval. The dashed diagonal line represents perfect calibration. WavefrontDiffusion exhibits curves that are consistently tighter to the diagonal, indicating superior calibration.
C.8 Decoding Dynamics Analysis
To explicitly address the concern regarding frontier evolution and stability, we visualize the step-wise decoding dynamics in Figure 4. This analysis captures the ”process” of generation, contrasting the rigid scheduling of BlockDiffusion with the adaptive flow of WavefrontDiffusion.
(a) Frontier Evolution
(b) Wavefront Confidence Heatmap
(c) Confidence Stability
Figure 4: Decoding Dynamics Analysis. (a) WavefrontDiffusion shows a smooth, continuous progression (green), avoiding the rigid ”staircase” artifacts of BlockDiffusion (red). (b) The heatmap reveals a clear ”diagonal frontier” where the wavefront adapts its velocity to token difficulty. (c) WavefrontDiffusion maintains consistently high post-finalization confidence, indicating minimal ”regret,” whereas BlockDiffusion suffers from confidence drops due to insufficient lookahead context.
1. Frontier Evolution (Smoothness vs. Stagnation). As shown in Figure 4(a), BlockDiffusion exhibits a characteristic ”staircase” pattern, where the frontier freezes within a block until all tokens are finalized, regardless of their difficulty. In contrast, WavefrontDiffusion (green curve) demonstrates a smooth, continuous progression. The absence of horizontal plateaus confirms that our method avoids premature freezing and adapts the generation pace naturally to the sequence structure.
2. The ”Wave” Mechanism (Adaptivity). Figure 4(b) visualizes the confidence heatmap of WavefrontDiffusion. The bright diagonal band represents the active wavefront. Notably, the boundary is not a perfect straight line but exhibits local fluctuations. This visualizes the adaptivity of the algorithm: the wavefront accelerates through high-confidence (easy) regions and slows down to accumulate context for low-confidence (hard) tokens, effectively implementing a dynamic computation allocation.
3. Post-Finalization Stability (Regret Analysis). In Figure 4(c), we track the average confidence of tokens after they are finalized. Ideally, a model should not ”regret” its decisions (i.e., confidence should not drop after finalizing). BlockDiffusion (red) shows jitter and drops, implying that forced block commits often lead to sub-optimal decisions that the model later finds inconsistent. WavefrontDiffusion (green) maintains a stable, near-1.0 confidence trajectory, proving that its context-aware expansion significantly reduces finalization regret.
Appendix D Theoretical Justification
In this section, we provide the formal theoretical framework that justifies WavefrontDiffusion as the optimal expansion of restricted-search decoding strategies. Our derivation rests on two core pillars: the Information Gradient Hypothesis and the optimality of dynamic restricted search.
D.1 Premise 1: The Information Gradient Hypothesis
Let $\mathcal{C}_{t}$ be the set of finalized tokens at step $t$. For any masked token $i$, let $d(i,\mathcal{C}_{t})$ be the minimum distance to the finalized context (as defined in Eq. 1). We introduce the Information Gradient Assumption, which states that the conditional entropy $H$ of a masked token $x_{i}$ is monotonically non-decreasing with respect to its distance from the finalized context:
$H(x_{i}|x_{\mathcal{C}_{t}})\leq H(x_{j}|x_{\mathcal{C}_{t}})\iff d(i,\mathcal{C}_{t})<d(j,\mathcal{C}_{t})$ (5)
This formalizes the intuition that tokens strictly adjacent to the known context possess the highest ”information density” and thus the highest probability of being correctly predicted.
D.2 Premise 2: The Validity of Restricted Search
The global optimization objective at step $t$ is to select a subset of tokens $S_{t}$ (where $|S_{t}|=K$) to finalize such that we maximize joint confidence. BlockDiffusion (Arriola et al., 2025) demonstrated a key theoretical contribution: Global search is unnecessary. It proved that limiting the search space to a constrained subset $\Omega\subset\mathcal{V}_{all}$ yields efficient results, provided the subset contains valid candidates.
  • BlockDiffusion’s Constraint: $\Omega_{Block}=\{i\mid i\in\text{Block}_{k}\}$.
  • Limitation: Because $\Omega_{Block}$ is static, there exists a non-zero probability of ”semantic mismatch,” where true high-confidence tokens lie outside the block. That is, $\exists j\notin\Omega_{Block}$ such that $H(x_{j})<\min_{i\in\Omega_{Block}}H(x_{i})$.
D.3 Theorem: WavefrontDiffusion as Optimal Restricted Search
Our method aligns the restriction boundary $\Omega$ with the Information Gradient defined in Premise 1. We define the Wavefront search space $\Omega_{Wave}$ as:
$\Omega_{Wave}=\{i\mid d(i,\mathcal{C}_{t})\leq R\}$ (6)
Theorem. Under the Information Gradient Assumption, the optimal subset $S^{*}$ (the set of $K$ tokens with lowest entropy) satisfies $S^{*}\subseteq\Omega_{Wave}$ for a sufficient radius $R$, whereas $S^{*}\subseteq\Omega_{Block}$ is not guaranteed.
Proof Sketch. Since entropy $H(x_{i})$ increases with distance $d(i,\mathcal{C}_{t})$, the ”lowest entropy isosurfaces” form expanding contours around $\mathcal{C}_{t}$:
  1. 1.
    BlockDiffusion forces a search within a rectangular window $\text{Block}_{k}$. This window inevitably intersects with high-entropy regions (far from context) while potentially excluding low-entropy regions (near context but in the next block) due to rigid segmentation.
  2. 2.
    WavefrontDiffusion strictly defines $\Omega_{Wave}$ to encapsulate the lowest distance $d$ values. Therefore, $\Omega_{Wave}$ is mathematically guaranteed to contain the highest density of correctable tokens.
  3. 3.
    Consequently, WavefrontDiffusion ensures that the local optimization is performed within the theoretical upper bound of candidate quality, extending the efficiency logic of BlockDiffusion to its optimal semantic conclusion.
Empirical Verification. This theoretical proof is empirically supported by our MHCO analysis in Section 4.4. The significantly lower MHCO scores for WavefrontDiffusion confirm that our dynamic boundary successfully captures the optimal local search space defined by this framework, whereas BlockDiffusion’s static boundary frequently violates it.